# On the Foundation of Distributionally Robust Reinforcement Learning

### v3

@@ 3.4 Satisfaction of the DPP: A Summary
    • A DPP always holds for the six cases outlined in Theorem 1, where neither the controller’s nor the adversary’s action set is assumed to be convex.
      An important scenario where the controller action distribution set becomes non-convex is when $\mathcal{Q}=\mathcal{Q}^{\mathrm{D}}$ is deterministic. In this context, Corollary 1.1 characterizes the solution of the Bellman equation in terms of the $q$-function.
    • Moreover, if the controller action distribution set $\mathcal{Q}$ is convex, then Theorem 4 implies the additional satisfaction of the DPP $\mu[u^{*}]=v^{*}(\mu,\Pi^{\mathrm{C}}_{\mathrm{H}},\mathrm{K}_{\mathrm{M}}^{\mathrm{S}})$ where the controller is history-dependent and the adversary is Markov.
    • When both the controller’s and the adversary’s action sets are convex and at least one is compact, the Sion’s minimax principle holds for the DR Bellman equation. In this case, the DPP invariably holds.
    • However, if the controller’s action set is non-convex, even with a convex adversary, the DPP generally holds only for the six cases specified in Theorem 1.
      Counterexamples for the other three cases can be found by considering $\mathcal{Q}=\mathcal{Q}^{\mathrm{D}}$, as detailed in Section 5.3.
@@
    • From the viewpoint of the interchange of sup-inf operations in the DR Bellman equation, if this interchange is valid, the DPP always holds. Moreover, the validity of this interchange at the level of the Bellman equation implies the interchange of the order of the two players in the dynamic decision-making environment defined by the value (2.16).
    • An important special case in which the interchange of sup-inf in the DR Bellman equation is consistently valid is when the adversary adopts an SA-rectangular strategy, and the controller is allowed to use deterministic policies, as shown in Theorem 3.
The two perspectives intersect when Sion’s minimax principle is satisfied by the DR Bellman equation. In such cases, both players have convex action sets, and the interchange of sup-inf operations is valid. Consequently, the DPP holds for all nine cases, see Corollary 2.1.
@@ 4 Optimality of Markov Time-Homogeneous Policies
A decision rule $\Delta=\left\{{\Delta(s)\in\mathcal{Q}:s\in S}\right\}$ is $\eta$-optimal for the Bellman equation (3.1) if for all $s\in S$
$u^{*}(s)\leq\Delta(s)[r(s,\cdot)]+\gamma\inf_{p_{s}\in\mathcal{P}_{s}}\Delta(s)\otimes p_{s}[u^{*}]+\eta.$
@@
Let $u^{*}$ be the solution to (3.1) and $\eta\geq 0$. Assume for some $\Pi=\Pi^{\mathrm{C}}_{\mathrm{H}},\Pi^{\mathrm{C}}_{\mathrm{M}},\Pi^{\mathrm{C}}_{\mathrm{S}}$ and $\mathrm{K}=\mathrm{K}_{\mathrm{H}}^{\mathrm{S}},\mathrm{K}_{\mathrm{M}}^{\mathrm{S}},\mathrm{K}_{\mathrm{S}}^{\mathrm{S}}$
$\mu[u^{*}]=v^{*}(\mu,\Pi,\mathrm{K}).$
Then, any $\eta$-optimal decision rule $\Delta$ for (3.1) induces a $\eta/(1-\gamma)$-optimal policy $\pi:=(\Delta,\Delta,\dots)\in\Pi^{\mathrm{C}}_{\mathrm{S}}$ for the maxmin control problem $v^{*}(\mu,\Pi,\mathrm{K})$ in the sense that for all $\mu\in\mathcal{P}(\mathcal{S})$
$0\leq v^{*}(\mu,\Pi,\mathrm{K})-\inf_{\kappa\in\mathrm{K}}v(\mu,\pi,\kappa)\leq\frac{\eta}{1-\gamma}.$
@@
Here, we intentionally allow $\eta=0$; i.e., if the assumptions in Theorem 5 hold and the solution of the Bellman equation induces a decision rule $\Delta=\left\{{\Delta(s):s\in S}\right\}$ that achieve the supremum in (3.1), then the Markov time-homogeneous policy $\pi:=(\Delta,\Delta,\dots)$ is optimal for the maxmin control problem.
Remark 5.
For the cases either $\mathcal{Q}=\mathcal{Q}^{\mathrm{D}}$ or the adversary policy class is SA-rectangular, Corollary 1.1 and Theorem 3 implies that any greedy deterministic decision rule
$d(s)\in\underset{a\in A}{\operatorname{arg}\,\operatorname{max}}\;q^{*}(s,a)$
is 00-optimal for the Bellman equation. Therefore, if the DPP holds (a green check in Table 4 or 1), by Theorem 5, the corresponding greedy policy $\pi:=(d,d,\dots)\in\Pi^{\mathrm{C}}_{\mathrm{S}}$ is optimal for the maxmin control problem.
The preceding remarks describe scenarios in which Markov time-homogeneous policies prove to be optimal. Additionally, Theorem 5 implies that Markov time-homogeneous policies, achieving arbitrarily small errors in terms of satisfying the DR Bellman equation (3.1), will yield near optimal performance with arbitrarily small optimality gap in the context of the maxmin control. Consequently, a DRRL procedure that identifies a policy approximating the solution of (3.1) is assured to achieve reasonable performance in addressing the maxmin control problem.
@@ 5 Absence of the DPP: Counterexamples
In this section, we provide counterexamples in the form of specific DRMDP instances for which the DPP does not hold. In particular, this section is organized as follows: In Section 5.1, we consider the setting where a history-dependent randomized controller is paired with a Markov time-homogeneous non-convex adversary, as indicated by the cross mark at (1,3)th entry in Table 3. In Section 5.2, we consider an instance with a Markov randomized controller and a Markov time-homogeneous non-convex adversary, represented by the cross mark at (2,3)th entry in Table 3. Finally, in Section 5.3, we establish counterexamples for all the cases with deterministic controllers in Table 4 for which a DPP is not proven.
@@ 5.1 History-dependent Randomized Controller Versus
We display the DRMDP instance used as the counterexample in Figure 1. This is used to verify the cross mark at (1,3)th entry in Table 3.
@@
The adversary’s action distribution set $\mathcal{P}_{I}:=\left\{{p_{I}^{(1)},p_{I}^{(2)}}\right\}$ in the $I$ state has two distributions characterized by
$p^{(1)}_{I,a_{1}}(B)=1\text{ and }p^{(1)}_{I,a_{2}}(G)=1,$
$p^{(2)}_{I,a_{1}}(G)=1\text{ and }p^{(2)}_{I,a_{2}}(B)=1.$
Furthermore, in states G,B, we both have
$p_{G}(I)=p_{B}(I)=1.$
We assume that reward function $r$ only depends on the states so that
$r(I)=0,\quad r(G)=1,\quad r(B)=-1.$
We consider a history-dependent policy $\pi$; At time 0, we randomly pick one action at state I. If we observe state G, we will choose the same action for the following time steps. If we observe state B, we will choose the alternative action for the following time steps.
For any Markov time-homogeneous adversary’s policy $\kappa$, we have
$v(I,\pi,\kappa)=0+0\times\gamma+0\times\gamma^{2}+\sum_{i=3,i\text{ odd}}^{\infty}\gamma^{i}=\frac{\gamma^{3}}{1-\gamma^{2}}>0.$ (5.1)
Here, for notation simplicity, we abuse the notation and write $v(I,\pi,\kappa):=v(\delta_{I},\pi,\kappa)$ as the value of state $I$ under the policy pair $(\pi,\kappa)$, where $\delta_{I}$ is the point mass measure at I. On the other hand, the DR Bellman equation corresponding to this instance is
$u(I)$ $=$ $0+\sup_{d\in\mathcal{P}(\mathcal{A})}\inf_{p_{I}\in\mathcal{P}_{I}}\gamma d\otimes p_{I}[u],$ (5.2)
$u(G)$ $=$ $1+\gamma u(I),$ (5.3)
$u(B)$ $=$ $-1+\gamma u(I).$ (5.4)
Solving the equations, we have that the solution $u^{*}$ to (5.2 - 5.4) is
$u^{*}(I)=0,\quad u^{*}(G)=1,\quad u^{*}(B)=-1.$
Notice that $u^{*}(I)<v(I,\pi,\kappa)$ in (5.1). By construction, we have $v^{*}(I,\Pi_{\mathrm{H}},\mathrm{K}_{\mathrm{S}}^{\mathrm{S}})\geq v(I,\pi,\kappa)>u^{*}(I)$. Therefore, the DPP is not satisfied.
In this case, due to the Markov time-homogeneity of the adversary, the task of the controller becomes a bandit learning problem (Lattimore and Szepesvári, 2020). Therefore, it is possible for the controller to implement a history-dependent policy that optimally “learns” the environment.
@@ 5.2 Markov Randomized Controller Versus
The adversary’s action distribution set $\mathcal{P}_{I}:=\left\{{p_{I}^{(1)},p_{I}^{(2)}}\right\}$ at state I contains two probability distributions:
$p^{(1)}_{I,a_{1}}$ $=$ $(1,0,0)^{\top}\quad\text{and}\quad p^{(1)}_{I,a_{2}}=\left(\frac{1}{2},\frac{1}{2},0\right)^{\top},$
$p^{(2)}_{I,a_{1}}$ $=$ $\left(0,\frac{1}{2},\frac{1}{2}\right)^{\top}\quad\text{and}\quad p^{(2)}_{I,a_{2}}=\left(\frac{1}{2},0,\frac{1}{2}\right)^{\top},$
where we use a column vector to denote the probabilities of states I,B,C.
Further, once the chain arrives state B or state C, it will stay in that state forever, i.e.,
$p_{B}(B)=p_{C}(C)=1.$
We assume again that rewards only depend on the states
$r(I)=r(C)=0,\quad r(B)=3/5$
and the discount factor $\gamma=0.8.$
We now proceed to solve the DR Bellman equation. It is easy to see that $u^{\ast}(C)=0,u^{\ast}(B)=3$. We claim that $u^{\ast}(I)=1$. By Proposition 1, the solution $u^{*}$ is unique, so if $u^{\ast}(C)=0,u^{\ast}(B)=3,u^{\ast}(I)=1$ verifies (3.1), it must be the solution. We proceed to check
$u^{\ast}(I)$ $=$ $\gamma\sup_{d\in\mathcal{P}(\mathcal{A})}\inf_{p_{I}\in\mathcal{P}_{I}}d(a_{1})p_{I,a_{1}}^{\top}u^{\ast}+d(a_{2})p_{I,a_{2}}^{\top}u^{\ast}$
$=$ $\gamma\sup_{d(a_{1})\in[0,1]}\inf\left\{d(a_{1})+\frac{1+3}{2}(1-d(a_{1})),\left(\frac{0+3}{2}\right)d(a_{1})+\frac{0+1}{2}(1-d(a_{1}))\right\}$
$=$ $0.8\sup_{d(a_{1})\in[0,1]}\inf\left\{2-d(a_{1}),\frac{1}{2}+d(a_{1})\right\}$
$=$ $1,$
where the optimal controller’s action distribution is $d(a_{1})=3/4.$
$\pi_{0}(I)=a_{1}\text{ and }\pi_{1}(I)=a_{2}\text{ and }\pi_{i}(I)=(3/4,1/4)\text{ for }i\geq 2.$
At time 0, if the adversary chooses $p_{I}^{(2)},$ we have
$v(I,\pi,p_{I}^{(2)})=0.8(u^{\ast}(B)+u^{\ast}(C))/2=1.2>1.$
On the other hand, if the adversary chooses $p_{I}^{(1)},$ we have
$v_{1}(I,\pi,p_{I}^{(1)})$ $=$ $\frac{\gamma}{2}(u^{\ast}(I)+u^{\ast}(B))=1.6,\text{ and}$
$v(I,\pi,p_{I}^{(1)})$ $=$ $\gamma v_{1}(A,\pi,p_{I}^{(1)})=1.44>1,$
where $v_{1}(I,\pi,p_{I}^{(1)})$ stands for the value of state A at time 1 under the Markov policy $\pi$ and the adversary with the choice $p_{I}^{(1)}$. Therefore, we conclude under the Markov non-time-homogeneous policy $\pi$, the controller can have a higher value than the solution of the DR Bellman equation, hence the absence of the DPP.
@@ 5.3 Deterministic Controllers
In this subsection, we focus on deterministic controllers. Specifically, we will use the DRMDP instance in Figure 3 to construct counterexamples to address the various cases with a red cross in Table 4. These cases include a history-dependent deterministic controller against a Markov convex adversary, a history-dependent deterministic controller paired with Markov time-homogeneous convex adversary, and a Markov deterministic controller with Markov time-homogeneous convex adversary.
I0I1I2IGB1/2121/21 / $21/$ 2
@@
We consider a DRMDP environment with six states I0,I1,I2,I,B,G. The initial state is always I0. At state I, there are two actions: ${A}=\{a_{1},a_{2}\}.$ For all other states, there is only one action. As shown in the Figure 3, at state I0, the transition probabilities are
$p_{I0}(I1)=p_{I0}(I2)=\frac{1}{2}.$
@@
The adversary has action distribution set $\mathcal{P}_{I}$ at state I. If we assume a convex adversary, we will use $\mathcal{P}_{I}:=\left\{{\alpha p_{I}^{(1)}+(1-\alpha)p_{I}^{(2)}:\alpha\in[0,1]}\right\}$, else if the adversary is assumed to be non-convex, then we define $\mathcal{P}_{I}:=\left\{{p_{I}^{(1)},p_{I}^{(2)}}\right\}$, where the two distributions $p_{I}^{(1)},p_{I}^{(2)}$ are characterized by
$p^{(1)}_{I,a_{1}}(B)=1\text{ and }p^{(1)}_{I,a_{2}}(G)=1,$
$p^{(2)}_{I,a_{1}}(G)=1\text{ and }p^{(2)}_{I,a_{2}}(B)=1.$
@@
In what follows, we only verify the case when the adversary is convex; i.e. $\mathcal{P}_{I}:=\left\{{\alpha p_{I}^{(1)}+(1-\alpha)p_{I}^{(2)}:\alpha\in[0,1]}\right\}$. The absence of the DPP is still valid when we assume a non-convex adversary with $\mathcal{P}_{I}:=\left\{{p_{I}^{(1)},p_{I}^{(2)}}\right\}$.
As before, we assume that rewards only depend on the states:
$r(I0)=0,r(I1)=0,r(I2)=0,r(I)=0,r(G)=1,r(B)=-1.$
We compute the solution $u^{*}(\cdot)$ of the DR Bellman equation:
$u(I)$ $=$ $0+\sup_{a\in\left\{a_{1},a_{2}\right\}}\inf_{p\in\mathcal{P}_{I}}\gamma p_{I,a}[u],$
$u(I0)$ $=$ $\frac{\gamma}{2}u(I1)+\frac{\gamma}{2}u(I2),u(I1)=\gamma u(I),u(I2)=\gamma u(I),$
$u(G)$ $=$ $1+\gamma u(I),u(B)=-1+\gamma u(I).$
The solution $u^{*}$ is
$u^{*}(I)=-\frac{\gamma}{1-\gamma^{2}},\quad u^{*}(I1)=u^{*}(I2)=-\frac{\gamma^{2}}{1-\gamma^{2}},$
$u^{*}(I0)=-\frac{\gamma^{3}}{1-\gamma^{2}},\quad u^{*}(G)=1-\frac{\gamma^{2}}{1-\gamma^{2}},\quad u^{*}(B)=-\frac{1}{1-\gamma^{2}}.$
@@
We first consider the case of a Markov controller with Markov time-homogeneous convex adversary. In particular, the controller uses the following Markov policy $\pi$ with
$\pi_{2}=a_{1},\quad\pi_{4}=a_{2},\quad\pi_{6}=a_{1},\quad\pi_{8}=a_{2},\ldots.$
We claim that this policy can achieve a higher reward than the solution of the DR Bellman equation. To see that, we use $p_{I}^{\alpha}$ to denote the adversary’s action $\alpha p_{I}^{(1)}+(1-\alpha)p_{I}^{(2)}$. Then, we have
$v_{2}(I,\pi,p_{I}^{\alpha})$ $=$ $0+\gamma\alpha v_{3}(B,\pi,p_{I}^{\alpha})+\gamma(1-\alpha)v_{3}(G,\pi,p_{I}^{\alpha}),$
$v_{3}(B,\pi,p_{I}^{\alpha})$ $=$ $-1+\gamma v_{4}(I,\pi,p_{I}^{\alpha}),v_{3}(G,\pi,p_{I}^{\alpha})=1+\gamma v_{4}(I,\pi,p_{I}^{\alpha}),$
$v_{4}(I,\pi,p_{I}^{\alpha})$ $=$ $0+\gamma\alpha v_{5}(G,\pi,p_{I}^{\alpha})+\gamma(1-\alpha)v_{5}(B,\pi,p_{I}^{\alpha}),$
$v_{5}(B,\pi,p_{I}^{\alpha})$ $=$ $-1+\gamma v_{6}(I,\pi,p_{I}^{\alpha}),v_{5}(G,\pi,p_{I}^{\alpha})=1+\gamma v_{6}(I,\pi,p_{I}^{\alpha}),$
$v_{6}(I,\pi,p_{I}^{\alpha})$ $=$ $v_{2}(I,\pi,p_{I}^{\alpha}).$
Solving the recursion, one obtains the solution and verifies the value $v(I0,\pi,p_{I}^{\alpha})>u^{\ast}(I0)$ uniformly in $\alpha$ as follows:
$v(I0,\pi,p_{I}^{\alpha})$ $=$ $(1-2\alpha)\frac{\gamma^{3}}{1+\gamma^{2}}\geq-\frac{\gamma^{3}}{1+\gamma^{2}}>u^{\ast}(I0),$
$v_{2}(I,\pi,p_{I}^{\alpha})$ $=$ $(1-2\alpha)\frac{\gamma}{1+\gamma^{2}}\geq-\frac{\gamma}{1+\gamma^{2}}>u^{\ast}(I),$
$v_{3}(B,\pi,p_{I}^{\alpha})$ $=$ $-1+\left(2\alpha-1\right)\frac{\gamma^{2}}{1+\gamma^{2}},v_{3}(G,\pi,p_{I}^{\alpha})=1+\left(2\alpha-1\right)\frac{\gamma^{2}}{1+\gamma^{2}}$
$v_{4}(I,\pi,p_{I}^{\alpha})$ $=$ $\left(2\alpha-1\right)\frac{\gamma}{1+\gamma^{2}}.$
@@
By Lemma 1, there must be a history-dependent deterministic controller that achieve no worse performance than $\pi$. So, the above constructions and inequalities also imply that a history-dependent deterministic controller with a Markov time-homogeneous convex adversary can achieve a higher value. Therefore, the DPP doesn’t hold for the two cases discussed above.

### v4

@@ 3.4 Satisfaction of the DPP: A Summary
    • A DPP always holds for the six cases in Theorem 1, where neither the controller’s nor the adversary’s action set is assumed to be convex. An important setting with a non-convex controller action distribution set is $\mathcal{Q}=\mathcal{Q}^{\mathrm{D}}$. In this context, Corollary 1.1 characterizes the solution of the Bellman equation using the $q$-function.
    • If the controller action distribution set $\mathcal{Q}$ is convex, then Theorem 4 implies the additional DPP $E_{\mu}[u^{*}(X_{0})]=v(\mu,\Pi_{\mathrm{H}},\mathrm{K}_{\mathrm{M}}^{\mathrm{S}})$ where the controller is history-dependent and the adversary is Markov.
    • When both the controller’s and the adversary’s action sets are convex and at least one is compact, Sion’s minimax principle holds for the robust Bellman equation. In this case, the DPP invariably holds.
    • However, if the controller’s action set is non-convex, even with a convex adversary, the DPP generally holds only for the six cases specified in Theorem 1. Counterexamples for the other three cases can be found by considering $\mathcal{Q}=\mathcal{Q}^{\mathrm{D}}$, as detailed in Section 5.3.
@@
    • If the robust Bellman equation shares the same solution as the interchanged equation (3.2), then the DPP holds, and the order of the two players can be interchanged in the dynamic decision-making environment characterized by the value (2.14).
    • An important special case where this interchange is always valid occurs when the adversary is SA-rectangular and the controller is permitted to use any deterministic policies, as established in Theorem 3.
The two perspectives intersect when Sion’s minimax principle is satisfied by the robust Bellman equation. In such cases, both players have convex action sets, and the interchange of sup-inf operations is valid. Consequently, the DPP holds for all nine cases, see Corollary 2.1.
@@ 4 Optimality of Markov Time-Homogeneous Policies
A decision rule $\Delta=\left\{{\Delta(\cdot|s)\in\mathcal{Q}:s\in S}\right\}$ is $\eta$-optimal for the Bellman equation (3.1) if for all $s\in S$, $u^{*}(s)\leq\inf_{p_{s}\in\mathcal{P}_{s}}E_{\Delta(\cdot|s),p_{s}}\left[r(s,A_{0})+\gamma u^{*}(X_{1})\right]+\eta.$
@@
Let $u^{*}$ be the solution to (3.1) and $\eta\geq 0$. Assume that $E_{\mu}[u^{*}(X_{0})]=v(\mu,\Pi,\mathrm{K})$ where $\Pi$ represents $\Pi_{\mathrm{H}}$, $\Pi_{\mathrm{M}}$, or $\Pi_{\mathrm{S}}$, and $\mathrm{K}$ corresponds to $\mathrm{K}_{\mathrm{H}}^{\mathrm{S}}$, $\mathrm{K}_{\mathrm{M}}^{\mathrm{S}}$, or $\mathrm{K}_{\mathrm{S}}^{\mathrm{S}}$. Then, any $\eta$-optimal decision rule $\Delta$ for (3.1) induces a $\eta/(1-\gamma)$-optimal policy $\pi:=(\Delta,\Delta,\dots)\in\Pi_{\mathrm{S}}$ for the maxmin control problem $v(\mu,\Pi,\mathrm{K})$ in the sense that for all $\mu\in\mathcal{P}(\mathcal{S})$, $0\leq v(\mu,\Pi,\mathrm{K})-\inf_{\kappa\in\mathrm{K}}v(\mu,\pi,\kappa)\leq\frac{\eta}{1-\gamma}.$
@@
Here, we intentionally allow $\eta=0$; i.e., if the assumptions in Theorem 5 hold and the solution of the Bellman equation induces a decision rule $\Delta=\left\{{\Delta(\cdot|s):s\in S}\right\}$ that achieve the supremum in (3.1), then the Markov time-homogeneous policy $\pi:=(\Delta,\Delta,\dots)$ is optimal for the maxmin control problem.
For the cases either $\mathcal{Q}=\mathcal{Q}^{\mathrm{D}}$ or the adversary policy class is SA-rectangular, Corollary 1.1 and Theorem 3 implies that any greedy deterministic decision rule $\pi(s)\in\underset{a\in A}{\operatorname{arg}\,\operatorname{max}}\;q^{*}(s,a)$ is 0-optimal for the Bellman equation. Therefore, if the DPP holds (a green check in Table 4 or 1), by Theorem 5, the corresponding greedy policy $(\pi,\pi,\dots)\in\Pi_{\mathrm{S}}$ is optimal for the maxmin control problem.
The preceding remarks clarify the settings in which Markov time-homogeneous policies are optimal. Furthermore, Theorem 5 implies that for RMDP models satisfying a DPP, policies from the Markov time-homogeneous class can achieve performance arbitrarily close to optimal. As a result, in these settings, a DRRL procedure that approximates the solution to (3.1) is guaranteed to deliver strong performance in solving the max-min control problem.
@@ 5 Absence of the DPP: Counterexamples
In this section, we provide counterexamples in the form of specific RMDP instances for which the DPP does not hold. The guiding insight is that DPP failure emerges precisely when the asymmetric information structures enable the controller to learn salient features of the worst-case adversary by experimenting over time. If the information structure lets the controller treat the interaction as a sequential exploration problem, much like a multi-armed bandit or RL task, then adaptive learning destroys the Markov optimality implied by a DPP. Viewing the controller through this “learning agent” lens proved crucial in designing the counterexamples that delineate the boundary between DPP validity and breakdown.
Concretely, this section is organized as follows: In Section 5.1, we consider the setting where a history-dependent randomized controller is paired with a Markov time-homogeneous non-convex adversary, as indicated by the cross mark at (1,3)th entry in Table 3. In Section 5.2, we consider an instance with a Markov randomized controller and a Markov time-homogeneous non-convex adversary, represented by the cross mark at (2,3)th entry in Table 3. Finally, in Section 5.3, we establish counterexamples for all the cases with deterministic controllers in Table 4 for which a DPP is not proven.
@@ 5.1 History-dependent Randomized Controller Versus
We display the RMDP instance used as the counterexample in Figure 1. This is used to verify the cross mark at (1,3)th entry in Table 3.
@@
[…]
Furthermore, in states G and B, the controller’s action set contains only one action, and the adversary’s uncertainty set includes only a single distribution, which always transitions back to the same state I. That is, $p_{\text{G}}(\text{I})=p_{\text{B}}(\text{I})=1.$ We assume that reward function $r$ only depends on the states so that $r(\text{I},a_{1})=r(\text{I},a_{2})=r(\text{I})=0,r(\text{G})=1,r(\text{B})=-1.$
We consider a history-dependent policy $\pi$; at time 0, we randomly pick one action at state I. If we observe state G, we will choose the same action for the following time steps. If we observe state B, we will choose the alternative action for the following time steps. We do not aim to prove that such a history-dependent policy $\pi$ is optimal; rather, we will only show that this policy achieves a higher value than what is implied by the DPP, which means the DPP is not satisfied.
For any Markov time-homogeneous adversary’s policy $\kappa$, we observe that at even time steps, the controller will always return to state I and receive a reward of 0. At odd time steps, the controller will receive a reward of either 1 or -1 with equal probability at step 1, and for steps $\geq 3$, the controller will always receive a reward of 1. Therefore, we have $v(\delta_{\text{I}},\pi,\kappa)=\sum_{i=3,i\text{ odd}}^{\infty}\gamma^{i}=\frac{\gamma^{3}}{1-\gamma^{2}}>0,$ where $\delta_{\text{I}}$ is the point mass measure at I. On the other hand, the robust Bellman equation corresponding to this instance is
$u(\text{I})=\sup_{\phi\in\mathcal{P}(\mathcal{A})}\inf_{p_{\text{I}}\in\mathcal{P}_{\text{I}}}\left(0+\gamma E_{\phi,p_{\text{I}}}[u(X_{1})]\right),\quad u(\text{G})=1+\gamma u(\text{I}),\quad u(\text{B})=-1+\gamma u(\text{I}).$ (5.1)
Solving the equations, we have that the solution $u^{*}$ to (5.1) is
$u^{*}(\text{I})=0,u^{*}(\text{G})=1,u^{*}(\text{B})=-1.$
Notice that $u^{*}(\text{I})<v(\delta_{\text{I}},\pi,\kappa)$. Therefore, the DPP is not satisfied.
In this case, due to the Markov time-homogeneity of the adversary, the task of the controller becomes a bandit learning problem [33]. Therefore, it is possible for the controller to implement a history-dependent policy that optimally “learns” the environment.
@@ 5.2 Markov Randomized Controller Versus
The adversary’s action distribution set $\mathcal{P}_{\text{I}}:=\left\{{p_{\text{I}}^{(1)},p_{\text{I}}^{(2)}}\right\}$ at state I contains two probability distributions:
$p^{(1)}_{\text{I},a_{1}}=(1,0,0),\quad p^{(1)}_{\text{I},a_{2}}=\left(\frac{1}{2},\frac{1}{2},0\right);\text{ and}\quad p^{(2)}_{\text{I},a_{1}}=\left(0,\frac{1}{2},\frac{1}{2}\right),\quad p^{(2)}_{\text{I},a_{2}}=\left(\frac{1}{2},0,\frac{1}{2}\right),$
where the transition probabilities to states I,B,C given the current state and action are represented as row vectors. Further, we notice that once the chain arrives state B or state C, it will stay in that state forever, i.e., $p_{\text{B}}(\text{B})=p_{\text{C}}(\text{C})=1.$
We again assume that rewards only depend on the states: $r(\text{I})=r(\text{C})=0,r(\text{B})=3/5$ and the discount factor $\gamma=0.8.$ In Lemma 2, we show that there exists a Markov non-time-homogeneous policy, the controller can have a higher value than the solution of the robust Bellman equation, hence the absence of the DPP. The proof of Lemma 2 is shown in Appendix D.
Lemma 2.
By solving the robust Bellman equation, we have $u^{\ast}(\mathrm{C})=0,u^{\ast}(\mathrm{B})=3,u^{\ast}(\mathrm{I})=1$. For a Markov but not time-homogeneous policy $\pi=(\pi_{0},\pi_{1},\dots)$ defined by $\pi_{0}(\mathrm{I})=a_{1}\text{ and }\pi_{1}(\mathrm{I})=a_{2}\text{ and }\pi_{i}(\mathrm{I})=(3/4,1/4)\text{ for }i\geq 2,$ we have $v(\delta_{\mathrm{I}},\pi,p_{\mathrm{I}}^{(1)})=1.44>1,\text{and }v(\delta_{\mathrm{I}},\pi,p_{\mathrm{I}}^{(2)})=1.2>1.$
@@ 5.3 Deterministic Controllers
In this subsection, we focus on deterministic controllers. Specifically, we will use the RMDP instance in Figure 3 to construct counterexamples to address the various cases with a red cross in Table 4. These cases include a history-dependent deterministic controller against a Markov convex adversary, a history-dependent deterministic controller paired with Markov time-homogeneous convex adversary, and a Markov deterministic controller with Markov time-homogeneous convex adversary.
$\text{I}_{0}$$\text{I}_{1}$$\text{I}_{2}$IGB1/21/21/21/2
@@
We consider an RMDP environment with six states $\text{I}_{0},\text{I}_{1},\text{I}_{2},\text{I}$,B,G. The initial state is always $\text{I}_{0}$. At state I, there are two actions: ${A}=\{a_{1},a_{2}\}.$ For all other states, there is only one action. As shown in the Figure 3, at state $\text{I}_{0}$, the transition probabilities are $p_{\text{I}_{0}}(\text{I}_{1})=p_{\text{I}_{0}}(\text{I}_{2})=\frac{1}{2}.$ As before, we assume that rewards only depend on the states: $r(\text{I}_{0})=0,r(\text{I}_{1})=0,r(\text{I}_{2})=0,r(\text{I})=0,r(\text{G})=1,r(\text{B})=-1.$
@@
The adversary has action distribution set $\mathcal{P}_{\text{I}}$ at state I. If we assume a convex adversary, we will use $\mathcal{P}_{\text{I}}:=\left\{{\alpha p_{\text{I}}^{(1)}+(1-\alpha)p_{\text{I}}^{(2)}:\alpha\in[0,1]}\right\}$, else if the adversary is assumed to be non-convex, then we define $\mathcal{P}_{\text{I}}:=\left\{{p_{\text{I}}^{(1)},p_{\text{I}}^{(2)}}\right\}$, where the two distributions $p_{\text{I}}^{(1)},p_{\text{I}}^{(2)}$ are characterized by
$p^{(1)}_{\text{I},a_{1}}(\text{B})=1,p^{(1)}_{\text{I},a_{2}}(\text{G})=1,\text{ and }p^{(2)}_{\text{I},a_{1}}(\text{G})=1,p^{(2)}_{\text{I},a_{2}}(\text{B})=1.$
@@
We will only verify the case when the adversary is convex; i.e. $\mathcal{P}_{\text{I}}:=\left\{{\alpha p_{\text{I}}^{(1)}+(1-\alpha)p_{\text{I}}^{(2)}:\alpha\in[0,1]}\right\}$. The absence of the DPP is still valid when we assume a non-convex adversary with $\mathcal{P}_{\text{I}}:=\left\{{p_{\text{I}}^{(1)},p_{\text{I}}^{(2)}}\right\}$.
Lemmas 3 and 4 demonstrate that a deterministic Markov controller can attain a value greater than that indicated by the robust Bellman equations. The proofs of these lemmas are provided in Appendix D.
Lemma 3.
By solving the robust Bellman equation, we have $u^{*}(\mathrm{I}_{0})=-\frac{\gamma^{3}}{1-\gamma^{2}}$.
@@
We first consider the case of a Markov controller with a Markov time-homogeneous convex adversary.
Lemma 4.
If controller uses the following Markov policy $\pi$ with $\pi_{2}=a_{1},\pi_{4}=a_{2},\pi_{6}=a_{1},\pi_{8}=a_{2},\ldots,$ we have $v(\delta_{\mathrm{I}_{0}},\pi,p_{\mathrm{I}}^{\alpha})=(1-2\alpha)\frac{\gamma^{3}}{1+\gamma^{2}}>u^{\ast}(\mathrm{I}_{0}).$
@@
By Lemma 1, there must be a history-dependent deterministic controller that achieves no worse performance than $\pi$. So, the above constructions and inequalities also imply that a history-dependent deterministic controller with a Markov time-homogeneous convex adversary can achieve a higher value. Therefore, the DPP doesn’t hold for the two cases discussed above.